June 29, 2016 Calibration

From GlueXWiki
Jump to: navigation, search

GlueX Calibration Meeting
Wednesday, June 29, 2016
11:00 am, EST
JLab: CEBAF Center, F326

Communication Information

Remote Connection

You can connect using BlueJeans using the meeting number 630 804 895 .       (Click "Expand" to the right for more details -->):

  1. Make sure you have created a BlueJeans account via your JLab CUE account using this link:

  2. Meeting ID: 630804895
    • (you may need to type this in, depending how you connect)

  3. If connecting via Web Browser: click this link (no passcode is needed):

  4. If connecting via iOS or Android App:
    • Use your JLab e-mail address to log in and then enter the meeting ID given above to join the meeting

  5. If connecting via Phone: Dial one of the following numbers and then enter the meeting ID above and hit "#" or "##"

  6. If connecting via Polycom unit:
    • Dial or bjn.vc
    • Enter meeting ID above
    • Use *4 to unmute


Talks can be deposited in the directory /group/halld/www/halldweb/html/talks/2016 on the JLab CUE. This directory is accessible from the web at https://halldweb.jlab.org/talks/2016/ .

Calibration Tasks

  1. Spring 2016 Recon Checklist
  2. Mcsmear updates - mcsmear 20160323
  3. Data Validation
  4. GlueX Detector Performance Document

Calibrations for Second Production Run

  • Fix handling of individually-read-out TAGM columns (Alex B.)
  • Tagger hit merging (Nathan/Alex B.)
  • Updates to FCAL non-linearity (Adesh)
  • Add timing information to FCAL clustering/inner ring calibrations (Adesh/Matt S.)
  • BCAL TDC calibrations (Andrei)
  • BCAL energy non-linearity correction (Will M.)
  • FDC wire timing & reconstruction (Lubomir/Simon/Alex A.)
  • TOF ADC energy and timing calibrations (Beni)
  • TOF alignment (Beni/Simon)
  • General timing realignment (Sean)
  • Updated error matrices (all)


  1. Announcements
  2. Calibration Updates
    • Calibrations for run group 4
    • Calibration Plans for Second Production Run
      • Calorimetry
        • FCAL (Adesh)
      • Tracking
      • TOF (Beni)
      • Start Counter
      • Photon Beam
  3. Simulations
    • sim1.1 for data/simulation matching studies (Sean/Mark I.)
    • mcsmear updates
  4. AOB


Attending: Sean (NU); Simon, Paul, Beni, Alex A., Alex B., Adesh, Justin, Mark I., Mark D., Eugene, Elton (JLab); Mike S., Naomi (CMU)

  1. Announcements
    • Mark Ito announced that /work/halld2 filled up this morning. The drive was filled with files from the recent launch. Alex A. is moving these files to /work/halld.
  2. We reviewed progress on calibrations for the last run group. The goal for these is to be finished by the end of Friday.
    • Alex B. is finishing up the TAGM calibrations and expects to be done by the end of today.
    • Adesh's FCAL calibration jobs are almost done, and he expects to be done by tomorrow.
    • Naomi is still working on CDC gain calibrations, but expects these to be easier than the 3rd run group. She reviewed her calibration work on those runs, which were early in the run period. The calibrations were complicated by the fact that the thresholds were set high, since the fADC algorithm was still being debugged. She found the temperature on the downstream side of the CDC changed by +-1 degree C, but the pressure varied between 98 and 102 kPa, sometimes quite quickly. A change in pressure of 1/10 of a kPa yielded a shift of ~2% in the MIP peak position. She took a set of good straws and used those to set the gains for these runs by scaling the gains from run 11059. Conditions in the next set of runs are more stable - for later runs she did find a linear relationship between pressure and gain correction, and is planning to make a parameterization that can be used for calibrating future runs.
    • Sean showed some distributions of tagger time alignment that illustrated the improvements in calibrations, and the variation in TAGM time resolution for different columns with the new gains. The difference in variation is due to increased noise in the low-yield fibers, which have higher gains applied..
  3. Calibration Progress
    • We reviewed the list of tasks planned to be finished for the next REST production run. Sean had added some items to the list based on discussions in other meetings. Mike S. pointed out that an important topic is to update the uncertainties being used e.g. for the kinematic fit and PID FOM calculations. Sean mentioned that these values need to be determined to update the simulations, and can be put in the reconstruction in parallel to that effort.
    • Adesh described his updates to the FCAL shower corrections in sim-recon. The previous parameterization combined a energy scale correction with the non-linear correction. This was needed since mcsmear writes out calibrated hits. Adesh has implemented some scale corrections to the hits in mcsmear. Matt Shepherd is working on the more comprehensive solution of having mcsmear write out digi (pre-calibrated) hits, similar to how the BCAL is simulated.
    • Beni has added an explicit ADC/TDC alignment step to his TOF calibrations and uploaded new values, so those should be better aligned. He also found a bug where timewalk corrections were not being correctly applied to the single-ended paddles. He is considering changing the timewalks corrections to use the ADC pulse amplitude rather than integral. This would involve some extensive changes and creation of new ROOT trees.
    • The next REST production run must be finished by the start of the next data taking period, so it should begin sometime in mid-to-late August.
  4. Simulations
    • Mark Ito has checked out the conditions files for sim1.1 and will start running test jobs.
    • He and Sean are working on a method to generate simulations with different run numbers and with event counts in proportion to the number of events in the real data. There are no run-dependent calibrations for simulations at the moment, but it was decided at the last Offline meeting to implement this now so that it is ready for when run-dependent calibrations for simulation are added.
    • It had been suggested to simulate events to total 10% of the data. This would be 2.6x10^9 events, which an order of magnitude larger than the sim1 run. A smaller amount, say 1-2% would be sufficient for studying data/MC matching.